# Gravitational Waves from First-Order Phase Transitions Assisted by Temperature-Enhanced Scatterings

### v1

@@ 3.2 Computation setup
To compute the stochastic GW spectrum generated by the first-order phase transition, we adopt the standard semi-analytic formalism developed in Refs. [31, 32, 36, 38]. In this framework, the total GW energy density spectrum today can be expressed as the sum of three main contributions:
$\Omega_{\mathrm{GW}}(f)h^{2}=\Omega_{\mathrm{col}}(f)h^{2}+\Omega_{\mathrm{sw}}(f)h^{2}+\Omega_{\mathrm{turb}}(f)h^{2},$ (3.1)
where the terms correspond respectively to GWs sourced by the collision of bubble walls ($\Omega_{\mathrm{col}}$), the long-lasting sound waves in the plasma ($\Omega_{\mathrm{sw}}$), and the turbulent motion of the plasma ($\Omega_{\mathrm{turb}}$) following the transition.
@@
$f_{\mathrm{peak}}\simeq 1.65\times 10^{-5}~{}\mathrm{Hz}\left(\frac{f_{*}}{\beta}\right)\left(\frac{\beta}{H}\right)\left(\frac{T_{n}}{100\,\mathrm{GeV}}\right)\left(\frac{g_{*}}{100}\right)^{1/6},$ (3.2)
where $T_{n}$ is the nucleation temperature, $\beta/H$ is the inverse duration parameter in units of the Hubble rate at $T_{n}$, $g_{*}$ is the number of relativistic degrees of freedom, and $f_{*}/\beta$ is an order-one numerical coefficient that depends on the detailed source dynamics.
@@
$\beta/H=T_{n}\left.\frac{d(S_{3}/T)}{dT}\right|_{T=T_{n}}.$ (3.3)
We use the package cosmoTransitions, [84], to solve the above equation. Once these thermal parameters are known, individual contributions $\Omega_{\mathrm{col}}$, $\Omega_{\mathrm{sw}}$, and $\Omega_{\mathrm{turb}}$ to the GW spectrum are evaluated using semi-analytic formulae that encode the known dependence on $\alpha$, $\beta/H$, $v_{w}$, and the nucleation temperature $T_{n}$. This approach captures the essential physics of GW generation from expanding and colliding bubbles, as well as the subsequent sound waves and turbulence in the plasma.
@@ 4 Parameter space analysis and phenomenology
$V_{\mathrm{eff}}(\phi,T)=V_{0}(\phi)+\Delta V_{\mathrm{thermal}}(\phi,T)+c\,T^{p}\phi^{2},$ (4.1)
@@
$p\in[-1.5,0],\quad c\in[0,c_{\max}],$ (4.2)
where $c_{\max}$ is chosen to preserve perturbativity and vacuum stability, we quantify how these parameters affect the dynamics of the electroweak-like phase transition and the resulting GW signatures.
@@ 4.1 Scanning the temperature-enhanced scattering parameters
$V_{\mathrm{eff}}(0,T_{c})=V_{\mathrm{eff}}(\phi_{\mathrm{min}}(T_{c}),T_{c}),$ (4.3)
@@
$\alpha$ $=\frac{\Delta\rho}{\rho_{\mathrm{rad}}(T_{n})},$ (4.5)
$\frac{\beta}{H}$ $=T_{n}\left.\frac{d}{dT}\left(\frac{S_{3}(T)}{T}\right)\right|_{T_{n}},$ (4.6)
@@
These trends show that the temperature-enhanced scattering term systematically enlarges the region of parameter space that supports strong, long-lasting first-order phase transitions.
@@ 4.2 Impact on phase transition parameters
$\alpha\uparrow\quad\Longrightarrow\quad\beta/H\downarrow,$ (4.7)
illustrated in Figure 5. The temperature-enhanced scattering term shifts and broadens this correlation, populating the region with $\alpha\gtrsim 0.1$ and $\beta/H\lesssim 100$, where GW production is most efficient.
@@ 4.3 Gravitational wave signal and detector prospects
$\Omega_{\mathrm{GW}}h^{2}(f)\approx\Omega_{\mathrm{peak}}\,S(f/f_{\mathrm{peak}}),$ (4.8)
@@
Overall, our analysis shows that temperature-enhanced scatterings significantly broaden the viable parameter space for strong first-order phase transitions and observable GWs, providing a robust, testable signature for future GW observatories.
@@ 5.1 Impact on phase transition strength
We begin by quantifying the effect of the temperature-enhanced scattering term (1.1) on the phase transition parameters: the strength parameter $\alpha$, inverse duration parameter $\beta/H$, critical temperature $T_{c}$, and nucleation temperature $T_{n}$.
By scanning over a range of exponents $p\in[-1.5,0]$ and coefficients $c\in[0,1]$, we numerically evaluate the temperature-dependent effective potential and compute the bounce action $S_{3}(T)/T$ as described in Section 2. The nucleation temperature $T_{n}$ is determined by the condition (4.4) ensuring that at least one bubble nucleates per Hubble volume. The strength parameter, $\alpha$ (4.5) quantifies the fraction of latent heat released relative to the background radiation energy density, while the inverse duration parameter, $\frac{\beta}{H}$ (4.6) measures how rapidly the transition completes.
@@ 5.2 Gravitational wave spectra
Next, we translate these phase transition parameters into predictions for the stochastic GW background. The peak amplitude of the GW energy density spectrum can be estimated as (4.8), where $S(f/f_{\mathrm{peak}})$ is a spectral shape function, and $\Omega_{\mathrm{peak}}$ depends on $\alpha$, $\beta/H$, $T_{n}$, and other efficiency factors, following the approach described in Section 3.
@@
  • By lowering $T_{n}$, they shift the peak frequency $f_{\mathrm{peak}}\propto T_{n}$ to lower values, closer to the sensitive bands of space-based interferometers.
@@ Appendix A Details of bounce action calculation
$\Gamma(T)\simeq T^{4}\left(\frac{S_{3}(T)}{2\pi T}\right)^{3/2}e^{-S_{3}(T)/T}.$ (A.1)
@@
$\frac{d^{2}\phi}{dr^{2}}+\frac{2}{r}\frac{d\phi}{dr}=\frac{dV_{\mathrm{eff}}(\phi,T)}{d\phi},$ (A.2)
@@
$\lim_{r\to\infty}\phi(r)=0,\quad\left.\frac{d\phi}{dr}\right|_{r=0}=0,$ (A.3)
@@
[…]

### v2

@@ 3.1 GW sources from FOPTs
For clarity, the total GW spectrum is expressed as a sum of three components,
$\Omega_{\mathrm{GW}}(f)h^{2}$ $=\Omega_{\mathrm{col}}(f)h^{2}+\Omega_{\mathrm{sw}}(f)h^{2}+\Omega_{\mathrm{turb}}(f)h^{2},$ (19)
with explicit forms given below [33, 40, 36]:
$\Omega_{\mathrm{col}}(f)h^{2}$ $=1.67\times 10^{-5}\left(\frac{H}{\beta}\right)^{2}\left(\frac{\kappa_{\phi}\,\alpha}{1+\alpha}\right)^{2}\left(\frac{100}{g_{*}}\right)^{1/3}\frac{0.11\,v_{w}^{3}}{0.42+v_{w}^{2}}S_{\mathrm{env}}\!\left(\frac{f}{f_{\mathrm{col}}}\right),$ (20)
$\Omega_{\mathrm{sw}}(f)h^{2}$ $=\Upsilon\cdot 2.65\times 10^{-6}\left(\frac{H}{\beta}\right)\left(\frac{\kappa_{\mathrm{sw}}\,\alpha}{1+\alpha}\right)^{2}\left(\frac{100}{g_{*}}\right)^{1/3}v_{w}\,S_{\mathrm{sw}}\!\left(\frac{f}{f_{\mathrm{sw}}}\right),$ (21)
$\Omega_{\mathrm{turb}}(f)h^{2}$ $={\cal O}(1)\times 3.35\times 10^{-4}\left(\frac{H}{\beta}\right)\left(\frac{\kappa_{\mathrm{turb}}\,\alpha}{1+\alpha}\right)^{3/2}\left(\frac{100}{g_{*}}\right)^{1/3}v_{w}\,S_{\mathrm{turb}}\!\left(\frac{f}{f_{\mathrm{turb}}}\right),$ (22)
where $S_{i}(f/f_{i})$ denote normalized spectral shapes for each source, and the efficiency factors $\kappa_{i}(\alpha,v_{w})$ represent the fraction of released vacuum energy converted into the corresponding channel. Among these, the sound-wave contribution typically dominates the integrated signal for subluminal wall velocities, while turbulence adds a high-frequency tail. We use the semi-analytic fits summarized above following the review [33] and[40] (bubble collisions), and the sound-shell model results of [36]. Recent numerical simulations indicate that the acoustic source can decay on timescales shorter than a Hubble time for sufficiently strong transitions; to account for this effect we multiply $\Omega_{\rm sw}$ by a suppression factor $\Upsilon$ (typically $\Upsilon=\min(1,H\tau_{\rm sw})$ or a fit derived from simulations). The turbulence contribution remains subject to larger modeling uncertainties and should be regarded as an order-of-magnitude estimate.
@@
The dominant contribution in our analysis comes from sound waves, for which the peak frequency in Eq. (8) corresponds to $f_{\mathrm{sw}}$. This is justified because the plasma motion persists for roughly a Hubble time after the transition, making the sound-wave source more efficient than the instantaneous envelope contribution from bubble collisions.
@@ 3.2 Computation setup
To compute the stochastic GW spectrum generated by the first-order phase transition, we adopt the standard semi-analytic formalism developed in Refs. [32, 33, 40, 42]. In this framework, the total GW energy density spectrum today can be expressed as the sum of three main contributions, as shown in Eq. 19.
@@
$f_{\mathrm{peak}}\simeq 1.65\times 10^{-5}~\mathrm{Hz}\left(\frac{f_{*}}{\beta}\right)\left(\frac{\beta}{H}\right)\left(\frac{T_{n}}{100\,\mathrm{GeV}}\right)\left(\frac{g_{*}}{100}\right)^{1/6}.$ (23)
This expression corresponds to the sound-wave source, with $f_{*}/\beta\simeq 0.9/(1.8-0.1v_{w}+v_{w}^{2})$ as given in [33]. For bubble collisions and turbulence, analogous relations hold with slightly shifted prefactors ($f_{*}/\beta\simeq 0.62/(1.8-0.1v_{w}+v_{w}^{2})$ for the envelope approximation and $f_{*}/\beta\simeq 1.4/(1+v_{w})$ for turbulence).
@@
$\beta/H=T_{n}\left.\frac{d(S_{3}/T)}{dT}\right|_{T=T_{n}}.$ (24)
All bounce actions and tunneling profiles were computed with the public package CosmoTransitions [95], into which we incorporated the $cT^{p}\phi^{2}$ term. The implementation modifies only the scalar mass matrix passed to CosmoTransitions; no additional thermal prefactors or mass resummations are introduced. For validation, we cross-checked representative points using an independent overshoot/undershoot routine described in Appendix A, finding agreement at the few-percent level.
The scattering-induced contribution $cT^{p}\phi^{2}$ is incorporated only through the field-dependent mass in the tree-level curvature entering $V_{\rm eff}(\phi,T)$. The remaining thermal free-energy contributions are computed using the exact bosonic and fermionic integrals $J_{B,F}(m^{2}/T^{2})$, without employing a high-temperature expansion. This avoids double counting: the same mediator loop responsible for $cT^{p}\phi^{2}$ is not reinserted into the thermal functions $J_{B,F}$, where the propagators are already thermally dressed. Consequently, the nucleation action and thermodynamic parameters $\{\alpha,\,\beta/H,\,T_{n}\}$ are evaluated from a potential in which the scattering correction appears once and only once, in agreement with the static-limit derivation of Appendix B.
Once these thermal parameters are known, individual contributions $\Omega_{\mathrm{col}}$, $\Omega_{\mathrm{sw}}$, and $\Omega_{\mathrm{turb}}$ to the GW spectrum are evaluated using the semi-analytic expressions above. This approach captures the essential physics of GW generation from expanding and colliding bubbles, as well as the subsequent sound waves and turbulence in the plasma.
The explicit spectral shapes and efficiency factors used in our computation are as follows:
[…]
The efficiency factors converting released vacuum energy into fluid motion are taken from [57]:
$\kappa_{\mathrm{sw}}(\alpha,v_{w})$ $\simeq\frac{\alpha}{0.73+0.083\sqrt{\alpha}+\alpha},$ (28)
$\kappa_{\phi}(\alpha,v_{w})$ $\simeq\frac{0.715\,\alpha+4\sqrt{3\alpha/2}}{1+0.715\,\alpha},$ (29)
$\kappa_{\mathrm{turb}}$ $=0.1\,\kappa_{\mathrm{sw}}.$ (30)
We include the acoustic suppression factor $\Upsilon$ [83],
$\Upsilon=1-\frac{1}{\sqrt{1+2\,H\,\tau_{\mathrm{sw}}}},\quad\tau_{\mathrm{sw}}\simeq\frac{R_{*}}{U_{f}},$ (31)
where $R_{*}$ is the mean bubble separation and $U_{f}$ is the root-mean-square fluid velocity. We apply $\Upsilon$ as a multiplicative damping of $\Omega_{\mathrm{sw}}$ to account for the finite lifetime of the acoustic source.
In all numerical plots, the total spectrum is the sum of the three components, $\Omega_{\mathrm{GW}}=\Omega_{\mathrm{col}}+\Omega_{\mathrm{sw}}+\Omega_{\mathrm{turb}}$, with sound waves dominating ($\gtrsim 90\%$ of the total signal), followed by bubble collisions ($\lesssim 10\%$) and turbulence ($\lesssim 5\%$). We take the turbulence prefactor $3.35\times 10^{-4}$ as calibrated in [33] and assume $v_{w}=0.9$ for the benchmark spectra unless stated otherwise. These prescriptions reproduce the spectral shapes observed in numerical simulations and ensure direct comparability with existing literature.
@@ 4 Parameter space analysis and phenomenology
$V_{\mathrm{eff}}(\phi,T)=V_{0}(\phi)+\Delta V_{\mathrm{thermal}}(\phi,T)+c\,T^{p}\phi^{2},$ (33)
@@
$p\in[-1.5,0],\quad c\in[0,c_{\max}],$ (34)
where $c_{\max}$ is chosen to preserve perturbativity and vacuum stability, we quantify how these parameters affect the dynamics of the electroweak-like phase transition and the resulting GW signatures. The upper limit $c_{\max}$ is fixed by requiring perturbativity and vacuum stability: (i) $|\Pi_{\phi}(T)|\lesssim 4\pi v^{2}$ for $T$ near the transition so that loop corrections remain perturbative, and (ii) the zero-temperature vacuum stays the global minimum. For the electroweak-like benchmarks considered, these conditions correspond to $c_{\max}\!\sim\!\mathcal{O}(0.1\!-\!1)$.
The effective scalar mass term in this setup is $m_{\mathrm{eff}}^{2}(T)=-\mu^{2}+a\,T^{2}+c\,T^{p}$, with $a>0$ the usual finite-temperature coefficient from one-loop self-energy diagrams and $cT^{p}$ the additional correction from temperature-enhanced scatterings. For $p<0$, the last term grows in relative importance as the Universe cools, steepening the curvature of the potential near the origin and thereby delaying the symmetry-breaking transition. Once the barrier separating the symmetric and broken minima disappears, the larger latent heat amplifies the strength parameter $\alpha$ and prolongs the duration of the transition, leading to smaller $\beta/H$.
@@ 4.1 Scanning the temperature-enhanced scattering parameters
$V_{\mathrm{eff}}(0,T_{c})=V_{\mathrm{eff}}(\phi_{\mathrm{min}}(T_{c}),T_{c}),$ (35)
@@
$\alpha$ $=\frac{\Delta\rho}{\rho_{\mathrm{rad}}(T_{n})},$ (37)
$\frac{\beta}{H}$ $=T_{n}\left.\frac{d}{dT}\left(\frac{S_{3}(T)}{T}\right)\right|_{T_{n}},$ (38)
@@
For numerical stability, we impose that the zero-temperature vacuum remains the global minimum and that $c\,T^{p}$ never drives $m_{\mathrm{eff}}^{2}(T)$ negative at high temperature. These criteria constrain $c_{\max}$ to the range $0.1\lesssim c_{\max}\lesssim 1$ for typical benchmark choices, ensuring the potential remains perturbative and stable.
@@
These behaviors follow directly from the modified thermal mass: a larger negative $p$ deepens the potential well more sharply near the transition, yielding stronger supercooling and a higher ratio of latent heat to background radiation. The smaller temperature derivative of the bounce action near $T_{n}$ reduces $\beta/H$, lengthening the GW-generating stage.
@@ 4.2 Impact on phase transition parameters
This delay arises because the additional term effectively adds a negative thermal correction to the curvature at the origin, flattening the potential and increasing the supercooling period. Consequently, bubbles nucleate only when the barrier becomes thin enough to overcome the thermal suppression, resulting in lower $T_{n}/T_{c}$ and larger $\alpha$.
@@
$\alpha\uparrow\quad\Longrightarrow\quad\beta/H\downarrow,$ (39)
illustrated in Figure 5. The temperature-enhanced scattering term shifts and broadens this correlation, populating the region with $\alpha\gtrsim 0.1$ and $\beta/H\lesssim 100$, where GW production is most efficient. This correlation provides an internally consistent cross-check: stronger supercooling (larger $\alpha$) implies a slower transition rate (smaller $\beta/H$), both consequences of the modified $m_{\mathrm{eff}}^{2}(T)$ profile. The result demonstrates that the temperature-dependent scattering term acts as a tunable handle on the phase transition dynamics without requiring fine-tuned tree-level parameters.
@@ 4.3 Gravitational wave signal and detector prospects
$\Omega_{\mathrm{GW}}h^{2}(f)\approx\Omega_{\mathrm{peak}}\,S(f/f_{\mathrm{peak}}),$ (40)
@@
Among these sources, the sound-wave contribution dominates in the parameter region identified here, since the bubble wall velocity remains subluminal ($v_{w}\lesssim 1$) and the released energy efficiently transfers into plasma motion rather than runaway walls. Consequently, the peak frequency in Eq. (40) corresponds to $f_{\mathrm{sw}}$, and the signal amplitude scales approximately as $\Omega_{\mathrm{GW}}\propto(\kappa_{\mathrm{sw}}\alpha)^{2}(\beta/H)^{-1}$.
@@
In summary, temperature-enhanced scatterings broaden the viable parameter space for strong first-order phase transitions by delaying nucleation, deepening the broken-phase minimum, and increasing latent heat release. The resulting decrease in $\beta/H$ and increase in $\alpha$ naturally boost the GW signal from sound waves, producing amplitudes that fall squarely within the detection reach of upcoming interferometers. This establishes a direct, testable link between microscopic temperature-dependent interactions and macroscopic cosmological observables.
@@ 5.1 Impact on phase transition strength
We begin by quantifying the effect of the temperature-enhanced scattering term (1) on the phase transition parameters: the strength parameter $\alpha$, inverse duration parameter $\beta/H$, critical temperature $T_{c}$, and nucleation temperature $T_{n}$.
By scanning over a range of exponents $p\in[-1.5,0]$ and coefficients $c\in[0,1]$, we numerically evaluate the temperature-dependent effective potential and compute the bounce action $S_{3}(T)/T$ as described in Section 2. The nucleation temperature $T_{n}$ is determined by the condition (36) ensuring that at least one bubble nucleates per Hubble volume. The strength parameter, $\alpha$ (37) quantifies the fraction of latent heat released relative to the background radiation energy density, while the inverse duration parameter, $\frac{\beta}{H}$ (38) measures how rapidly the transition completes.
@@
These trends follow directly from the temperature-dependent effective mass, $m_{\mathrm{eff}}^{2}(T)=-\mu^{2}+aT^{2}+cT^{p}$. For $p<0$, the additional $cT^{p}$ term effectively delays the restoration of symmetry by maintaining a positive curvature at the origin until lower $T$. As a result, the potential barrier persists longer, leading to increased supercooling before bubble nucleation. When tunneling finally occurs, the larger vacuum energy difference enhances the latent heat (larger $\alpha$), while the gentler slope of $S_{3}(T)/T$ around $T_{n}$ suppresses $\beta/H$. Both effects strengthen the GW source.
@@ 5.2 Gravitational wave spectra
Next, we translate these phase transition parameters into predictions for the stochastic GW background. The peak amplitude of the GW energy density spectrum can be estimated as (40), where $S(f/f_{\mathrm{peak}})$ is a spectral shape function, and $\Omega_{\mathrm{peak}}$ depends on $\alpha$, $\beta/H$, $T_{n}$, and other efficiency factors, following the approach described in Section 3.
@@
  • By lowering $T_{n}$, they shift the peak frequency $f_{\mathrm{peak}}\propto T_{n}$ to lower values, bringing the spectrum into the optimal sensitivity range of space-based interferometers.
@@
Quantitatively, for sound-wave–dominated transitions the GW peak scales as
$\Omega_{\mathrm{GW}}^{\text{peak}}\!\propto\!\left(\frac{H}{\beta}\right)\!\left(\frac{\kappa_{\mathrm{sw}}\alpha}{1+\alpha}\right)^{\!2},\quad f_{\mathrm{peak}}\!\propto\!\frac{\beta}{H}\frac{T_{n}}{100~\mathrm{GeV}}.$
Hence, the reduction in $\beta/H$ and $T_{n}$ achieved for negative $p$ simultaneously enhances $\Omega_{\mathrm{GW}}$ and shifts $f_{\mathrm{peak}}$ into the LISA/DECIGO/BBO bands. We have verified numerically that $\Omega_{\mathrm{sw}}$ dominates over $\Omega_{\mathrm{col}}$ and $\Omega_{\mathrm{turb}}$ for the range $0.6\!\lesssim\!v_{w}\!\lesssim\!1$, consistent with the expectation of plasma-dominated FOPTs.
@@
We find that parameter choices around $(p,c)\!\sim\!(-1,0.3)$ typically lead to $\alpha\!\sim\!0.1$$0.3$, $\beta/H\!\sim\!50$$100$, and $T_{n}/T_{c}\!\sim\!0.7$, corresponding to $\Omega_{\mathrm{GW}}^{\text{peak}}h^{2}\!\sim\!10^{-12}$$10^{-10}$, well within the projected sensitivity of LISA and DECIGO. Stronger scatterings ($c\!\gtrsim\!1$) can further enhance $\alpha$ but risk destabilizing the potential or exceeding perturbativity limits.
@@
In summary, the inclusion of the $cT^{p}\phi^{2}$ term not only modifies the thermal mass of the order parameter but also directly controls the macroscopic observables $(\alpha,\beta/H,T_{n})$ that determine GW production. This mechanism provides a continuous bridge between microphysical temperature-dependent scattering processes and the observable GW spectrum, thereby offering a novel and testable probe of beyond–Standard Model dynamics in the early Universe.
@@ 6 Discussion and Conclusion
From a thermal field-theoretic perspective, this term should be viewed as an effective parametrization of the temperature-dependent self-energy corrections induced by scatterings whose amplitudes scale as $\langle\sigma v\rangle\!\propto\!T^{p}$. In the Matsubara formalism, the finite-temperature propagator acquires thermal mass corrections $\Pi(T)\!\sim\!\!\int d^{3}k\,n_{B}(k)\,\sigma(T)$, and when $\sigma(T)$ increases at low $T$, the corresponding contribution to the scalar two-point function behaves as $\delta m_{\phi}^{2}(T)\!\propto\!T^{p}$ with $p<0$. The resulting quadratic operator $cT^{p}\phi^{2}$ therefore encapsulates the leading effect of these enhanced scatterings on the thermal potential, while higher-order or nonlocal terms are parametrically suppressed. This construction ensures that the inclusion of temperature-dependent scattering effects in $V_{\mathrm{eff}}$ is fully consistent with finite-temperature quantum field theory and provides a compact, model-independent way to capture the relevant dynamics.
@@
Finally, the phenomenological parametrization introduced here offers a bridge between microscopic scattering physics and macroscopic cosmological observables. Detecting a stochastic GW signal consistent with a temperature-enhanced FOPT would provide indirect evidence for temperature-dependent self-energy effects or light-mediator interactions in the early Universe, thereby linking cosmological GW measurements with laboratory-scale particle dynamics.
@@ Appendix A Details of bounce action calculation
$\Gamma(T)\simeq T^{4}\left(\frac{S_{3}(T)}{2\pi T}\right)^{3/2}e^{-S_{3}(T)/T}.$ (41)
@@
$\frac{d^{2}\phi}{dr^{2}}+\frac{2}{r}\frac{d\phi}{dr}=\frac{dV_{\mathrm{eff}}(\phi,T)}{d\phi},$ (42)
@@
$\lim_{r\to\infty}\phi(r)=0,\quad\left.\frac{d\phi}{dr}\right|_{r=0}=0,$ (43)
@@
All bounce profiles and actions reported in the main text were obtained using the public package CosmoTransitions [95], into which the temperature-enhanced term $c\,T^{p}\phi^{2}$ was added. To verify numerical stability and to cross-check the package output, we also implemented an independent overshoot/undershoot solver following the procedure outlined below. The two methods agree at the few-percent level across representative benchmark points, confirming that the results shown in the figures are robust.
@@
$S_{3}(T)=4\pi\int_{0}^{\infty}dr\,r^{2}\left[\frac{1}{2}\left(\frac{d\phi}{dr}\right)^{2}+V_{\mathrm{eff}}(\phi,T)\right].$ (44)
@@
For each temperature point, the bounce equation is solved until both the profile and its derivative vanish within relative tolerance $10^{-5}$ at the outer boundary. We verify that the difference between the CosmoTransitions and our independent implementation does not exceed a few percent for all benchmark parameters, ensuring the reliability of the extracted nucleation temperatures and transition parameters used in the main analysis.